OpenStudy (love_333):

Medal and fan! Angle a: 8x - 2 Angle b: 2x - 8 Angle C: 94 - 4x List the sides of triangle ABC in order from shortest to longest.

OpenStudy (love_333):

A: AC AB BC B: AC BC AB C: AB AC BC D: BC AC AB

OpenStudy (love_333):

@StudyGurl14 @Ms-Brains @boldjon @Brill @jhonyy9 anyone?

OpenStudy (anonymous):

What I would do is use draw a graph and find the points of intersection for the triangles Then using those points and the distance formula, you can do this problem

jhonyy9 (jhonyy9):

so you know that opposite a small angle is a short side - yes ? so just give value for example x=1 calculi the value of angles and check to the opposite side what value you have got - ok. ?

jhonyy9 (jhonyy9):

so you know that opposite to angle a will be the side BC and opposite to angle b will be AC and opposite to angle c will be AB - try calculi for x=1 what value will get for angle a ?

OpenStudy (love_333):

If you did x = 1 you get 6 for A

jhonyy9 (jhonyy9):

sorry you need giving for example 5 - for the second not give you minus - yes ?

jhonyy9 (jhonyy9):

so than Angle a: 8x - 2 Angle b: 2x - 8 Angle C: 94 - 4x than you give value x=5 will get angle a=8*5-2 = 40-2 =38 angle b=2*5-8=10-8=2 angle c=94-4*5=94-20=74 so and now because you need make the list from shortest to longest than see the smaller angle is angle b so opposite to angle b is AC so result the side AC will be the first in this list ok. ?

jhonyy9 (jhonyy9):

do you can continue it now ?